OpenStudy (en):

What mass could be lifted by the centrifugal force of an object having a mass of 98g if it were revolving in a horizontal circle, 20cm in diameter, and making 240 revolutions/min?

OpenStudy (mww):

Try using \[F = \frac {mv ^{2}}{r} \] Convert revolutions per minute into metres per second.

OpenStudy (mrnood):

this question is flawed in many ways. a) It is centriPETAL force that causes circular motion (centrFUGAL force is not a real force) b) centripetal force does no work c) The question says the force is horizontal - therefroe it cannot 'lift' anything. d) WHY use grams , cm - go with kg and m to get with the metric system IF it meansd what could the EQUIVALENT magnitude of the force lift then use the equation above to work out F then - the weight of the mass is Mg so Mg = mv^2/r (M is mass lifted, m= 98 grams)